Circular No. 6859

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998T IN NGC 3690 H. Yamaoka, Kyushu University; T. Kato, Kyoto University; A. V.

Filippenko, University of California at Berkeley; and S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology, report: “The position of SN 1998T reported on IAUC 6844 is incorrect; the position derived from an R-band image taken at the 0.95-m Ayabe Astronomical Observatory reflector on Mar. 18.5 UT (by M. Yamamoto) is α = 11^{h}28^{m}33^{s}.16, δ = +58^{o}33′43′′.7 (equinox 2000.0; based on GSC positions). Comparison with the nearby star (R = 17.3, IAUC 6844) on the Ayabe image shows that SN 1998T was not brighter than R = 15.5, and was possibly at R > 16.0, much dimmer than the reported visual magnitude estimates around this epoch (see below). This discrepancy could be due to contamination by the host galaxy or to misidentification: some galaxy catalogues, including RC3, incorrectly identify IC 694; the designation of Wynn-Williams et al. (see IAUC 6830) is also incorrect, and it was the basis for originally assigning SN 1998T to IC 694. NGC 3690 (Markarian 171) is an interacting pair of galaxies, and IC 694 is a small E/S0 galaxy located 1′ northwest of NGC 3690; these three compose Arp 299. Note that NGC 3690 produced several supernovae in the past decade: SN 1992bu (IAUC 5960), SN 1993G (IAUC 5718), and the probable radio supernova reported on IAUC 4988. Keck-2 spectra of the brightest knots, obtained by Filippenko on Mar. 27, show that SN 1998T is the only clear supernova in NGC 3690 at this time. A chart of SN 1998T is available at

http://www.obs.misato.wakayama.jp/~ayabe/michi/ayabe.jpg.”

D. Balam reports that a check of his CCD image (cf. IAUC 6844) shows no point source at the position specified above; he suggests that SN 1998T may be hidden in the scatter caused by the galaxy core on his frame. K. Hornoch and M. Plsek, Lelekovice, Czech Republic, reported on Mar. 28 that observations made on Mar. 26 show nothing at the correct position of SN 1998T down to visual mag 16, and conclude that their earlier observations (cf. IAUC 6843, 6844) were not of SN 1998T and should therefore be disregarded.

COMET 69P/TAYLOR Visual m_{1} estimates: Feb. 14.83 UT, 11.3 (A. Kammerer, Ettlingen,

Germany, 0.20-m reflector); Mar. 1.99, 11.9 (J. Warell, Sandvreten, Sweden, 0.38-m refl.); 20.95, 11.8 (K. Sárneczky, Ágasvár, Hungary, 0.44-m refl.); 30.82, 12.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1998 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green